Take a paragraph of about 100 words in English from any of your textbook . Frequency distribution table of the vowelsvowels paragraph and also draw a bar graph frequency distribution table.


1
Expert's answer
2021-07-02T07:44:59-0400

We can calculate the frequency of the each vowel of the text using the next formula:


"f(v)=n(v)/n(all)*100\% ,"

where f(v) - frequency of the certain vowel, n(v) - amount of its occurences in this text, n(all) - the number of the all symbols of the text. So we have


"f(A)=n(A)/n(all)=46/600=7.67\%"

"f(E)=n(A)/n(all)=63/600=10.5\%"

"f(I)=n(I)/n(all)=34/600=5.67\%"

"f(O)=n(O)/n(all)=61/600=10.17\%"

"f(U)=n(U)/n(all)=19/600=3.17\%"

"f(Y)=n(Y)/n(all)=13/600=2.17\%"

We can unite all the results in the frequency table and build the bar of its distribution: